A father wants to have R16000 available for his sons education on his 18th birthday. Calculate the amount he must invest on his sons 10th birthday at 14% compounded interest?

1
Expert's answer
2021-11-16T11:21:38-0500

No of years "18-10=8 years"


Rate = "\frac{14}{100}=0.14"


"16000=P(1+0.14)^8"


"16000=P(1.14)^8"


"P=\frac{16000}{(1.14)^8}=5608.94"


The amount he should invest is R5608.94